9. 3% H2O2 contains 30.0 g of H2O2 per litre. If 5.00 mL of H2O2 is decomposed, how many moles of oxygen gas will be produced?
1
Expert's answer
2020-04-28T12:21:24-0400

2 H2O2 = 2 H2O + O2

5 mL contains 5*30/1000 = 0.15 g H2O2 which is 0.15 g / 34 g/mole = 4.4*10^(-3) mole of H2O2.

There will be 2.2*10^(-3) mole of O2 oxygen gas